OpenStudy (anonymous):

Integrate: sin^3(x) dx

OpenStudy (anonymous):

= 3sin²x * cosx

OpenStudy (anonymous):

First, we should break-up the sin^3(x) because that isn't useful at all: \[\int\limits_{}^{} \sin^3(x) dx = \int\limits_{}^{} \sin^2(x) \sin(x) dx\] By the Pythagorean Identity, we know Sin^2(x) = 1 - Cos^2(x), so: \[\int\limits_{}^{} Sin^2(x) Sin(x) dx = \int\limits_{}^{} (1 - Cos^2x) Sin(x) dx\] now, do you have any ideas of what you should do? Hmm... let's go ahead and try a u-substitution: u = Cos(x) du = - Sin(x) dx - du = Sin(x) dx \[\int\limits_{}^{} (1 - Cos^2x) Sin(x) dx = \int\limits_{}^{} (1 - u^2)(-du)\] \[= -\int\limits\limits_{}^{} (1 - u^2)(du)\] now lets integrate that to get: \[- (u - (1/3)u^3)\] Now back substituting: \[= - (Cos(x) - (1/3)Cos^3 x)\] \[= - Cos(x) + (1/3)Cos^3 (x)\]

OpenStudy (anonymous):

thank you!